On an x-y coordinate plane, suppose x represents the year that a car was sold and y represents the sale price of the car. Rochelle bought a used car for $14,000 in 2005. She sold the car for $8,000 in 2010. Use the slope of the line connecting the points representing the two car sales to calculate the depreciation of the car's value per year over the time Rochelle owned it. (Enter the depreciation as a positive value.)
__________ dollars per year

Answer by stanbon(75887)

On an x-y coordinate plane, suppose x represents the year that a car was sold and y represents the sale price of the car.
Rochelle bought a used car for $14,000 in 2005.
She sold the car for $8,000 in 2010. Use the slope of the line connecting the points representing the two car sales to calculate the depreciation of the car's value per year over the time Rochelle owned it.
------
You have two points::(5,14000) and (10,8000)
------
slope = (8000-14000)/(10-5) = -6000/5 = -1200
------
(Enter the depreciation as a positive value.)
$1200 dollars per year
--------------
Cheers,
Stan H.
------------